OpenStudy (jazzyfa30):

Britney is preparing boxes of assorted chocolates. Chocolate covered peanuts cost $7 per pound. Chocolate covered caramel calls $6.50 per pound. The boxes of assorted candies contain five more pounds of peanut candies than caramel candies. if the total amount sold was $575, and many pounds of each candy were needed to make the boxes? Select one: a. 25 lbs of peanut candies and 40 lbs of carmel candies b. not enough information given: no solution c. 35 lbs of peanut candies and 30 lbs of carmel candies d. 45 lbs of peanut candies and 40 lbs of carmel candies

OpenStudy (jazzyfa30):

@nicolebears can you help e with this

OpenStudy (whpalmer4):

let P be the amount of chocolate covered peanuts, and C be the amount of chocolate covered caramels. P costs $7/lb, so the value of the peanuts sold is 7P. Similarly, C costs $6.50/lb, so the value of the caramels sold is 6.5P. We also know that each box contains 5 pounds more peanuts than caramels. Overall, that means we'll sell more peanuts than caramels, but how much? At least we can immediately strike one of the answer choices from consideration: a) has more caramels than peanuts, so it would be difficult to do that if each box contains more peanuts than caramels. Next, we can figure out the value of each combination, and see if it totals $575. If it does not, that answer is not a valid choice.

OpenStudy (jazzyfa30):

so B is out to cus it says no solution ????

OpenStudy (jazzyfa30):

i think the answers c ????

OpenStudy (jhannybean):

Try creating 2 equations in which you can figure out how many peanuts and how many caramels.

OpenStudy (jazzyfa30):

i did that

OpenStudy (jhannybean):

Which equations did you come up with? can you write them up here? :)

OpenStudy (jazzyfa30):

I did 6.5c*x=575 and 7p*x=575 ???

OpenStudy (jazzyfa30):

is it right @whpalmer4

OpenStudy (whpalmer4):

no, you have both caramels and peanuts being sold, so those equations are not correct. remember that P = C + 5 because 5 lbs more peanuts are sold than caramels.

OpenStudy (jazzyfa30):

so d

OpenStudy (whpalmer4):

well, check your work: 7P + 6.5C = 575 P = 45 C=40 7(45)+6.5(40) = 315 + 260 = 575 so that answer satisfies both equations (because 40+5=45)

OpenStudy (jazzyfa30):

ok i got the answer